OpenStudy (anonymous):

9.01] What is the value of the coefficient "a" when the quadratic equation y = (3x − 4)(2x − 1) is written in standard form? 6 5 −11 −4 I think it is 5, am I right?

OpenStudy (anonymous):

not familiar with the coefficient "a" if you could explain it i can give you an answer

OpenStudy (anonymous):

y = (3x − 4)(2x − 1) = 6x^2-11x+4 = 6(x^2 - 11x/6 + 1/24) = 6(x-11/12)^2 +... now, u can identify with standard form of qudratic equation : y=a(x-p)^2 + q

OpenStudy (anonymous):

It is the variable that comes before x in the quadratic equations. :)

OpenStudy (anonymous):

Is it 6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, thx. :) :)

OpenStudy (anonymous):

welcome :)